OpenStudy (anonymous):

Solve. 1/5(25-5a)=4-a A. –1 B. 0 C. no solution D. all real numbers

OpenStudy (ccswims):

Well first using the distributive property, do 1/5 * 25 and 1/5 * 5a

OpenStudy (ccswims):

Can you do that for now?

OpenStudy (anonymous):

k let me see, im really bad at these

OpenStudy (anonymous):

would it be 1/5 * 5-5a= 4-a?

OpenStudy (ccswims):

No, first do 1 divided by 5, that gives you 0.2. Now multiply 0.2 * 25 and 0.2 * 5a. What do you get?

OpenStudy (anonymous):

5

OpenStudy (ccswims):

good now whats 0.2 * 5?

OpenStudy (anonymous):

1

OpenStudy (ccswims):

yes, and 1a can be simplified to just "a". So now you have 5 - a = 4 - a Now you can add/subtract "like terms". So cancel out 5 and do 4 - 5

OpenStudy (anonymous):

-1

OpenStudy (ccswims):

And now add a+a (2a) So we end up with -1 = 2a To get "a" by itself we cancel out 2 and divide -1 by 2

OpenStudy (anonymous):

-0.5=a

OpenStudy (ccswims):

yes, and if we plug that in for "a" we see it is not correct. therefore the answer is C "no solution". To check even further (I did this just to make sure, but you can too) try plugging in any number above 0 or equal to 0, you will also see that both sides do not even out

OpenStudy (anonymous):

thnx! you really helped me!